# Efficient Multimodal Learning from Data-centric Perspective

### v2

@@ 2.1 Large Language Model
Recent advancements in Large Language Model (LLM) such as GPT [18, 19, 3], LLaMA [20, 7] and PaLM [21, 8] have demonstrated remarkable capacity in various natural language processing tasks, for example, dialogue, creative writing, and problem-solving. Benefiting from scalable Transformer-based architecture [6] design and web-scale training data source, these models have become foundation models for general reasoning tasks. Among many research endeavors, one line of work focuses on miniaturizing LLM, for more efficient and low-cost training and inference. By leveraging high-quality pre-training and instruction tuning data, along with optimized lightweight architecture, lightweight LLMs with fewer than 3 billion parameters, such as Phi [12, 13], StableLM-2 [22], and TinyLLaMA [23], have demonstrated performance comparable to larger models with 7/13 billion parameters.
@@ 2.2 Multimodal Large Language Model
Following the success of LLM, one line of research recently shifted further toward the Multimodal Large Language Model (MLLM) [3, 1] for unified cross-modality understanding and reasoning. Building upon various pre-trained LLM backbones, various methods are proposed to perform cross-modality fusion. Flamingo [2] and BLIP-2 [24] propose different modality fusion techniques to fuse visual tokens to frozen large language models through gated attention or query transformers. Inspired by the success of instruction tuning, LLaVA [25, 26] and MiniGPT-4 [27, 28] use visual instruction tuning to align visual input to LLMs, demonstrating remarkable success. More recent work, such as Kosmos-2 [4] and Shikra [5], further empowers MLLM with grounded visual understanding capacity. While these models exhibit promising potential for general-purpose visual reasoning and planning tasks, they are, in general, extremely expensive and prohibitive to train and deploy.
@@ 3.2 Training Data Construction
We construct a high-quality pre-training data from LAION-2B [34], i.e., we condense LAION-2B into a 2M coreset for data-efficient learning. Specifically, we design a three-step coreset selection scheme based on CLIP embedding. Firstly, motivated by SemDeDup [35], we cluster all 2B image embeddings by k-means, and then in each cluster, build an undirected graph that any two embeddings are connected if their cosine similarity is above a predetermined threshold. Only one sample is kept in each connected sub-graph whose Euclidean distance to the cluster centroid ranks at the median. By setting the threshold to , we obtain a subset of 952M. Secondly, we sort the remaining samples by the cosine similarity between their text embedding and image embedding and keep samples ranking , subsequently resulting in a subset of 190M. In this way, the low-quality image-text pairs are removed. Thirdly, we sort the remaining samples by the cosine similarity between its image embedding and its cluster centroid and keep samples ranking , resulting in a subset of 38M, capturing the essence and diversity of LAION-2B. We finally randomly sampled 2 million samples from the 38M corset, resulting in Bunny-pretrain-LAION-2M for an appropriate training cost.
@@
We collect a set of visual instruction tuning datasets – DataOptim111https://github.com/BAAI-DCAI/DataOptim, based on which we explore the better combination of fine-tuning datasets. Specifically, we leverage SVIT-mix-665K [17] and replace ShareGPT-40K [26] in it with WizardLM-evol-instruct-70K [36], resulting to Bunny-695K. We find that tuning MLLM on multimodal data may damage its cognition ability inherited from the pre-trained LLM. Probably, it is caused by the less informative and less diverse text in the multimodal training data. Keeping several high-quality pure text data in the fine-tuning dataset can relieve this problem. In experiments, we jointly use Bunny-695K and LLaVA-665K in the finetuning stage for better performance.
@@ 3.3 Training Procedure
We employ a two-stage training strategy. The first stage is the pre-training stage, where we align the visual embeddings from a pre-trained vision encoder with the text embeddings from the LLM. The second stage is the fine-tuning stage, where we apply visual instruction tuning to fully utilize the capabilities of the MLLM across various multimodal tasks. In both stages, we use the same cross-entropy loss for next-token prediction. During the pre-training stage, only the cross-modality projector is optimized for one epoch. In the fine-tuning stage, we use LoRA [37] to train the vision encoder, the cross-modality projector, and the LLM backbone for one epoch. For more training details, please refer to our GitHub repository. We observe that LoRA empirically leads to better performance compared to fully tuning all components of the model. This improvement is likely because smaller models are more prone to catastrophic forgetting, which LoRA tuning helps to mitigate.
@@ 4.1 Settings
We evaluate Bunny on eleven popular benchmarks: MME perception [38], MME cognition [38], MMBench [39] test split, MMBench [39] dev split, SEED-Bench-1 [40], MMMU [41] validation split, MMMU [41] test split, VQA-v2 [42] test-dev split, GQA [43] test-dev-balanced split, ScienceQA-IMG [44] test split and POPE [45]: the averaged F1-score of three categories (random, popular and adversarial) on the validation set of MSCOCO, to throughout assess its multimodal capacity. We compare against state-of-the-art multimodal large language models (MLLMs), including BLIP-2 [24], InstructBLIP [46], Shikra-13/7B [5], SVIT-v1.5-3B [17], LVIS-INSTRUCT4V-13B [47], VILA1.5-13B/3B [48], Mini-Gemini-HD-13B [49], LLaVA-v1.5-13/7B [26], LLaVA-Next-13B [50], MiniGPT-v2 [28], mPLUG-Owl2 [51], SPHINX-Intern2 [52], Yi-VL-6B [53], DeepSeek-VL-7B [54], MM1-7B/3B-Chat [55], Idefics2[56], LLaVA-Next-Llama-3-8B [57], TinyGPT-V [29], MC-LLaVA [58], Moondream2 [59], MobileVLM-V2-3B [60], TinyLLaVA-share-Sig-Phi [61], Mipha-3B [62], MiniCPM-V-2 [16], Imp-v1.5-4B-Phi3 [10].
@@ 4.2 Comparison to the State of the Art
| TinyGPT-V [29] | – | – | – | – | – | – | – | – | 33.6 | – | – |
| MC-LLaVA [58] | – | – | – | – | – | – | – | 64.2 | 49.6 | – | 80.6 |
| Moondream2 [59] | – | – | – | – | – | – | – | 79.4 | 63.1 | – | – |
| MobileVLM-V2-3B [60] | 1440.5 | – | – | 63.2 | – | – | – | – | 61.1 | 70.0 | 84.7 |
| TinyLLaVA-share-Sig-Phi [61] | 1464.9 | – | – | 66.9 | – | – | – | 79.9 | 62.0 | 69.1 | 86.4 |
| VILA1.5-3B [48] | 1442.4 | – | – | 63.4 | – | – | – | 80.4 | 61.5 | 69.0 | 85.9 |
| Mipha-3B [62] | 1488.9 | 295.0 | – | 69.7 | – | – | – | 81.3 | 63.9 | 70.9 | 86.7 |
| MiniCPM-V-2 [16] | 1411.4 | 396.8 | – | 69.6 | – | 38.2 | – | – | – | – | – |
| Imp-v1.5-4B-Phi3 [10] | 1507.7 | – | – | 73.3 | – | – | – | 81.5 | 63.5 | 78.0 | 86.9 |
@@
| Bunny-4B | 1503.9 | 362.9 | 74.1 | 74.1 | 64.6/71.7 | 40.2 | 38.8 | 81.7 | 63.4 | 76.3 | 87.0 |
As shown in Tab. 1, Bunny-4B and Bunny-8B outperform MLLMs with similar sizes on most of the benchmarks, demonstrating exceptional multimodal understanding and reasoning capabilities. Remarkably, Bunny-8B achieves an 80.4 score on ScienceQA-IMG, which overwhelms the runner-up MM1-7B-Chat (72.6) by a large margin. Bunny-4B also shows superiority over MM1-3B-Chat on the challenging MMMU benchmark, namely 6.3% and 5.1% improvements on the validation and test splits respectively.
@@ 4.3 Ablation Study
| MME | MME | MMB | MMB | SEED | MMMU | MMMU | VQA | GQA | SQA | POPE | |
| Fully Tuning | 1465.2 | 277.5 | 67.9 | 67.2 | 62.1 | 36.8 | 33.1 | 79.1 | 61.7 | 72.8 | 85.9 |
| LoRA | 1488.8 | 289.3 | 69.2 | 68.6 | 62.5 | 38.2 | 33.0 | 79.8 | 62.5 | 70.9 | 86.8 |
@@
As mentioned before, we find that LoRA empirically leads to better performance than fully tuning. The possible reason is that fully tuning on the small number of SFT data may lead to the catastrophic forgetting. We show the experimental results in Tab. 2.
| Fine-tuning Data | Epoch | MME | MME | MMB | MMB | SEED | MMMU | MMMU | VQA | GQA | SQA | POPE |
| Bunny-695K | 1 | 1488.8 | 289.3 | 69.2 | 68.6 | 62.5 | 38.2 | 33.0 | 79.8 | 62.5 | 70.9 | 86.8 |
| Bunny-695K | 2 | 1480.6 | 272.5 | 71.2 | 69.0 | 63.0 | 37.6 | 32.5 | 81.0 | 64.0 | 69.7 | 86.3 |
| Bunny-695K + LLaVA-665K | 1 | 1501.3 | 272.5 | 72.4 | 71.1 | 63.0 | 37.8 | 33.4 | 81.0 | 63.6 | 69.5 | 87.2 |
@@
Tab. 3 shows the ablation study on data strategy. Compared to the baselines – training on Bunny-695K for one and two epochs respectively, training on Bunny-695K + LLaVA-665K for one epoch can generally bring performance improvements on multiple benchmarks. The results indicate that increasing the data diversity can promote performance under the same training cost.
| Vision Encoder | MME | MME | MMB | MMB | SEED | MMMU | MMMU | VQA | GQA | SQA | POPE | |
| Bunny-695K + LLaVA-665K (LoRA) | Frozen | 1501.3 | 272.5 | 72.4 | 71.1 | 63.0 | 37.8 | 33.4 | 81.0 | 63.6 | 69.5 | 87.2 |
| Bunny-695K + LLaVA-665K (LoRA) | Trainable | 1521.5 | 291.8 | 71.9 | 71.3 | 63.1 | 38.0 | 33.1 | 81.1 | 63.7 | 70.8 | 87.0 |
@@
As shown in Tab. 4, enabling the vision encoder trainable in the fine-tuning stage is better than freezing its pre-trained parameters.
@@
| Vision Encoder | LLM | MME | MME | MMB | MMB | SEED | MMMU | MMMU | VQA | GQA | SQA | POPE |
| EVA02-CLIP-L (0.4B) | Phi-1.5 (1.3B) | 1213.7 | 278.9 | 60.9 | 56.8 | 56.4 | 30.0 | 28.4 | 76.5 | 60.4 | 58.2 | 86.1 |
| StableLM-2 (1.6B) | 1301.0 | 235.0 | 58.4 | 56.4 | 55.3 | 29.8 | 29.4 | 74.6 | 56.7 | 60.0 | 84.8 | |
| Phi-2 (2.7B) | 1421.0 | 285.4 | 68.6 | 67.4 | 62.2 | 35.9 | 32.6 | 78.9 | 62.3 | 69.1 | 87.1 | |
| SigLIP-SO (0.4B) | Phi-1.5 (1.3B) | 1230.0 | 237.5 | 61.2 | 59.7 | 57.7 | 30.0 | 29.1 | 78.0 | 61.1 | 61.3 | 85.8 |
| StableLM-2 (1.6B) | 1366.8 | 236.1 | 65.1 | 62.8 | 58.8 | 29.9 | 29.8 | 78.9 | 60.9 | 61.1 | 85.9 | |
| Phi-2 (2.7B) | 1488.8 | 289.3 | 69.2 | 68.6 | 62.5 | 38.2 | 33.0 | 79.8 | 62.5 | 70.9 | 86.8 | |
| Phi-3-Mini (3.8B) | 1503.9 | 362.9 | 74.1 | 74.1 | 71.7 | 40.2 | 38.8 | 81.7 | 63.4 | 76.3 | 87.0 | |
| Llama-3-8B | 1607.8 | 324.6 | 76.5 | 76.3 | 73.6 | 43.7 | 38.8 | 82.3 | 64.5 | 80.4 | 87.4 |

### v3

@@ 2.1 Large Language Model
Recent advancements in Large Language Model (LLM) such as GPT [18, 19, 3], LLaMA [20, 7, 21] and PaLM [8, 22] have demonstrated remarkable capacity in various natural language processing tasks, for example, dialogue, creative writing, and problem-solving. Benefiting from scalable Transformer-based architecture [6] design and web-scale training data source, these models have become foundation models for general reasoning tasks. Among many research endeavors, one line of work focuses on miniaturizing LLM, for more efficient and low-cost training and inference. By leveraging high-quality pre-training and instruction tuning data, along with optimized lightweight architecture, lightweight LLMs with fewer than 3 billion parameters, such as Phi [12, 13], StableLM-2 [23], and TinyLLaMA [24], have demonstrated performance comparable to larger models with 7/13 billion parameters.
@@ 2.2 Multimodal Large Language Model
Following the success of LLM, one line of research recently shifted further toward the Multimodal Large Language Model (MLLM) [3, 1] for unified cross-modality understanding and reasoning. Building upon various pre-trained LLM backbones, various methods are proposed to perform cross-modality fusion. Flamingo [2] and BLIP-2 [25] propose different modality fusion techniques to fuse visual tokens to frozen large language models through gated attention or query transformers. Inspired by the success of instruction tuning, LLaVA [26, 27, 28, 29] and MiniGPT-4 [30, 31] use visual instruction tuning to align visual input to LLMs, demonstrating remarkable success. More recent work, such as Kosmos-2 [4] and Shikra [5], further empowers MLLM with grounded visual understanding capacity. While these models exhibit promising potential for general-purpose visual reasoning and planning tasks, they are, in general, extremely expensive and prohibitive to train and deploy.
@@ 3.2 Training Data Construction
We construct a high-quality pre-training data from LAION-2B [36], i.e., we condense LAION-2B into a 2M coreset for data-efficient learning. Specifically, we design a three-step coreset selection scheme111https://github.com/BAAI-DCAI/Dataset-Pruning/tree/main/LAION based on CLIP embedding. Firstly, motivated by SemDeDup [37], we cluster all 2B image embeddings by k-means, and then in each cluster, build an undirected graph that any two embeddings are connected if their cosine similarity is above a predetermined threshold. Only one sample is kept in each connected sub-graph whose Euclidean distance to the cluster centroid ranks at the median. By setting the threshold to , we obtain a subset of 952M. Secondly, we sort the remaining samples by the cosine similarity between their text embedding and image embedding and keep samples ranking , subsequently resulting in a subset of 190M. In this way, the low-quality image-text pairs are removed. Thirdly, we sort the remaining samples by the cosine similarity between its image embedding and its cluster centroid and keep samples ranking , resulting in a subset of 38M, capturing the essence and diversity of LAION-2B. We finally randomly sampled 2 million samples from the 38M corset, resulting in Bunny-pretrain-LAION-2M for an appropriate training cost.
@@
We collect a set of visual instruction tuning datasets – DataOptim222https://github.com/BAAI-DCAI/DataOptim, based on which we explore the better combination of fine-tuning datasets. Specifically, we leverage SVIT-mix-665K [17] and replace ShareGPT-40K [27] in it with WizardLM-evol-instruct-70K [38], resulting to Bunny-695K. We find that tuning MLLM on multimodal data may damage its cognition ability inherited from the pre-trained LLM. Probably, it is caused by the less informative and less diverse text in the multimodal training data. Keeping several high-quality pure text data in the fine-tuning dataset can relieve this problem. In experiments, we jointly use Bunny-695K, LLaVA-665K [27] and ALLaVA-Instruct-4V [39] in the fine-tuning stage for better performance.
@@ 3.3 Training Procedure
We employ a two-stage training strategy. The first stage is the pre-training stage, where we align the visual embeddings from a pre-trained vision encoder with the text embeddings from the LLM. The second stage is the fine-tuning stage, where we apply visual instruction tuning to fully utilize the capabilities of the MLLM across various multimodal tasks. In both stages, we use the same cross-entropy loss for next-token prediction. During the pre-training stage, only the cross-modality projector is optimized for one epoch. In the fine-tuning stage, we generally use LoRA [40] to train the LLM backbone, and the vision encoder and the cross-modality projector are fully tuned. In sec. 4.3, we conduct ablation studies about LoRA, data, epoch and freezing component, and we also explore high image resolution and weight merging for more practical MLLMs. For more training details, please refer to our GitHub repository.
@@ 4.1 Settings
We evaluate Bunny on eleven popular benchmarks: MME perception [41], MME cognition [41], MMBench [42] test split, MMBench [42] dev split, SEED-Bench-1 [43], MMMU [44] validation split, MMMU [44] test split, VQA-v2 [45] test-dev split, GQA [46] test-dev-balanced split, ScienceQA-IMG [47] test split and POPE [48]: the averaged F1-score of three categories (random, popular and adversarial) on the validation set of MSCOCO, to throughout assess its multimodal capacity. We compare against state-of-the-art multimodal large language models (MLLMs), including GPT-4V [49], BLIP-2 [25], InstructBLIP [50], Shikra-13/7B [5], SVIT-v1.5-13B [17], LVIS-INSTRUCT4V-13B [51], ShareGPT4V-13B [52], VILA1.5-13/3B [53], Mini-Gemini-HD-13B [54], LLaVA-v1.5-13/7B [27], LLaVA-NeXT-13/7B [28], MiniGPT-v2 [31], mPLUG-Owl2 [55], SPHINX-Intern2 [56], Yi-VL-6B [57], DeepSeek-VL-7B [58], MM1-7/3B-Chat [59], Idefics2 [60], LLaVA-NeXT-Llama-3-8B [29], MobileVLM-V2-3B [61], TinyLLaVA-Phi-2-SigLIP-3.1B [62], Mipha-3B [63], MiniCPM-V 2.0 [64], Imp-v1.5-4B-Phi3 [10].
@@ 4.2 Comparison to the State of the Art
MobileVLM-V2-3B [61] | 1440.5 | – | – | 63.2 | –/– | – | – | – | 61.1 | 70.0 | 84.7 |
TinyLLaVA-Phi-2-SigLIP-3.1B [62] | 1466.4 | – | – | – | –/– | 38.4 | – | 80.1 | 62.1 | 73.0 | 87.2 |
VILA1.5-3B [53] | 1442.4 | – | – | 63.4 | 60.9/67.9 | 33.3 | 30.8 | 80.4 | 61.5 | 69.0 | 85.9 |
Mipha-3B [63] | 1488.9 | 295.0 | – | 69.7 | –/68.9 | – | – | 81.3 | 63.9 | 70.9 | 86.7 |
MiniCPM-V 2.0 [64] | 1411.4 | 396.8 | 69.1 | 69.6 | –/67.1 | 38.2 | – | – | – | 80.7 | 86.3 |
@@
Imp-v1.5-4B-Phi3 [10] | 1507.7 | – | – | 73.3 | –/– | – | – | 81.5 | 63.5 | 78.3 | 86.9 |
Bunny-4B | 1581.5 | 361.1 | 75.7 | 74.2 | 64.9/72.5 | 41.4 | 38.4 | 82.1 | 63.2 | 78.3 | 87.2 |
As shown in Tab. 1, Bunny-4B (SigLIP-SO + Phi-3-Mini) and Bunny-8B (SigLIP-SO + Llama-3-8B) outperform MLLMs with similar sizes on most of the benchmarks, demonstrating exceptional multimodal understanding and reasoning capabilities. Remarkably, Bunny-8B achieves an 1644.1 score on MME, which overwhelms the runner-up LLaVA-NeXT-Llama-3-8B (1603.7) by a large margin. Bunny-4B also shows superiority over MM1-3B-Chat on the challenging MMMU benchmark, namely 7.5 and 4.7 improvements on the validation and test splits respectively. Besides, Bunny-4B and Bunny-8B accept images up to 11521152.
@@ 4.3 Ablation Study
LLM Tuning | MME | MME | MMB | MMB | SEED(-I) | MMMU | MMMU | VQA | GQA | SQA | POPE |
Fully Tuning | 1465.2 | 277.5 | 67.9 | 67.2 | 62.1(69.5) | 36.8 | 33.1 | 79.1 | 61.7 | 72.8 | 85.9 |
LoRA | 1488.8 | 289.3 | 69.2 | 68.6 | 62.5(70.6) | 38.2 | 33.0 | 79.8 | 62.5 | 70.9 | 86.8 |
@@
In the fine-tuning stage, we observe that LoRA empirically leads to better performance compared to fully tuning the LLM. This improvement is likely because smaller models are more prone to catastrophic forgetting, which LoRA tuning helps to mitigate. We show the experimental results in Tab. 2.
For below experiments, the LLM is always tuned using LoRA.
Fine-tuning Data | Epoch | MME | MME | MMB | MMB | SEED(-I) | MMMU | MMMU | VQA | GQA | SQA | POPE |
Bunny-695K | 1 | 1488.8 | 289.3 | 69.2 | 68.6 | 62.5(70.6) | 38.2 | 33.0 | 79.8 | 62.5 | 70.9 | 86.8 |
Bunny-695K | 2 | 1480.6 | 272.5 | 71.2 | 69.0 | 63.0(71.3) | 37.6 | 32.5 | 81.0 | 64.0 | 69.7 | 86.3 |
Bunny-695K + LLaVA-665K | 1 | 1501.3 | 272.5 | 72.4 | 71.1 | 63.0(71.1) | 37.8 | 33.4 | 81.0 | 63.6 | 69.5 | 87.2 |
Fine-tuning Data | Epoch | MME | MME | MMB | MMB | SEED(-I) | MMMU | MMMU | VQA | GQA | SQA | POPE |
Bunny-695K | 1 | 1614.7 | 310.0 | 75.6 | 75.4 | 65.2(72.6) | 42.3 | 38.2 | 81.5 | 63.4 | 80.4 | 87.3 |
Bunny-695K + ALLaVA-Instruct-4V | 1 | 1599.6 | 348.9 | 76.3 | 75.6 | 66.2(73.4) | 41.6 | 38.3 | 81.9 | 64.1 | 79.5 | 86.4 |
@@
Tab. 3 shows the ablation study on data strategy. Jointly leveraging Bunny-695K and LLaVA-665K for one epoch generally yields performance improvements across multiple benchmarks. A comparison between training on Bunny-695K for two epochs and training on Bunny-695K and LLaVA-665K for one epoch demonstrates that increasing data diversity enhances performance under the same training cost. To further enhance diversity, we incorporate ALLaVA-Instruct-4V [39]. Given the detailed answers provided in ALLaVA-Instruct-4V, we append “Please answer the question in detail.” to the end of questions to avoid overfitting the model to long-form answers. As shown in Tab. 4, adding ALLaVA-Instruct-4V brings performance improvements. Based on above results, we set the fine-tuning stage to be 1 epoch in the rest experiments in this paper.
Fine-tuning Data | Vision Encoder | MME | MME | MMB | MMB | SEED(-I) | MMMU | MMMU | VQA | GQA | SQA | POPE |
Bunny-695K | Frozen | 1488.8 | 289.3 | 69.2 | 68.6 | 62.5(70.6) | 38.2 | 33.0 | 79.8 | 62.5 | 70.9 | 86.8 |
Trainable | 1486.0 | 281.1 | 69.5 | 69.3 | 62.3(70.6) | 37.7 | 33.2 | 79.9 | 62.6 | 70.6 | 85.9 | |
Bunny-695K + LLaVA-665K | Frozen | 1501.3 | 272.5 | 72.4 | 71.1 | 63.0(71.1) | 37.8 | 33.4 | 81.0 | 63.6 | 69.5 | 87.2 |
Trainable | 1521.5 | 291.8 | 71.9 | 71.3 | 63.1(71.1) | 38.0 | 33.1 | 81.1 | 63.7 | 70.8 | 87.0 |
LLM | Fine-tuning Data | MME | MME | MMB | MMB | SEED(-I) | MMMU | MMMU | VQA | GQA | SQA | POPE |
Phi-3-Mini | B-695K + L-665K | 1503.5 | 336.4 | 74.4 | 74.3 | 64.2(71.4) | 38.8 | 38.6 | 81.6 | 63.1 | 75.0 | 87.0 |
B-695K + L-665K + A-Ins | 1590.1 | 343.9 | 75.6 | 74.4 | 64.4(71.9) | 42.4 | 38.1 | 82.0 | 62.9 | 77.8 | 86.9 | |
Llama-3-8B | B-695K + L-665K | 1571.2 | 314.3 | 75.9 | 75.5 | 64.8(72.6) | 42.2 | 37.6 | 82.4 | 64.8 | 79.2 | 86.8 |
B-695K + L-665K + A-Ins | 1649.7 | 341.1 | 77.2 | 76.5 | 65.5(73.1) | 40.4 | 38.8 | 82.8 | 64.1 | 79.0 | 87.3 |
@@
By default, we freeze the vision encoder in both stages of our primary experiments. We also investigate the influence of enabling the vision encoder to be trainable during the fine-tuning stage. As shown in Tab. 5, the results vary depending on the fine-tuning data used. When employing Bunny-695K + LLaVA-665K, unfreezing the vision encoder leads to better results. Tab. 6 shows that the performance can be further enhanced with additional finetuning data when the vision encoder is trainable.
Pre-training LR | MME | MME | MMB | MMB | SEED(-I) | MMMU | MMMU | VQA | GQA | SQA | POPE |
1488.8 | 289.3 | 69.2 | 68.6 | 62.5(70.6) | 38.2 | 33.0 | 79.8 | 62.5 | 70.9 | 86.8 | |
1494.3 | 277.5 | 69.6 | 69.0 | 62.1(70.5) | 36.4 | 33.5 | 80.0 | 62.3 | 72.9 | 86.9 | |
1448.3 | 261.1 | 70.5 | 68.8 | 62.4(70.4) | 36.8 | 33.3 | 80.0 | 62.3 | 71.1 | 86.4 |
LoRA LR | Vision & Projector LR | MME | MME | MMB | MMB | SEED(-I) | MMMU | MMMU | VQA | GQA | SQA | POPE |
1521.5 | 291.8 | 71.9 | 71.3 | 63.1(71.1) | 38.0 | 33.1 | 81.1 | 63.7 | 70.8 | 87.0 | ||
1492.8 | 283.9 | 71.2 | 69.7 | 63.1(70.9) | 38.4 | 33.3 | 80.4 | 62.8 | 71.5 | 86.8 | ||
987.7 | 240.4 | 0.0 | 0.0 | 25.9(26.3) | 27.9 | 25.3 | 69.0 | 53.7 | 36.7 | 83.3 |
Learning Rate.
As shown in Tab. 7, increasing the pre-training learning rate does not consistently lead to improvements. In this context, both and are relatively good choices, while is sub-optimal. As shown in Tab. 8, the original learning rates set in LLaVA are the most appropriate. A learning rate that is too high causes extremely inferior results, including several sudden spikes in loss during training.
@@
Fine-tuning Data | Vision Encoder | Res | MME | MME | MMB | MMB | SEED(-I) | MMMU | MMMU | VQA | GQA | SQA | POPE |
Bunny-695K + LLaVA-665K | Trainable | 384 | 1521.5 | 291.8 | 71.9 | 71.3 | 63.1(71.1) | 38.0 | 33.1 | 81.1 | 63.7 | 70.8 | 87.0 |
448 | 1489.2 | 309.6 | 71.6 | 71.5 | 63.3(71.4) | 37.2 | 32.5 | 81.4 | 63.2 | 71.0 | 87.2 |
Fine-tuning Data | Vision Encoder | Res | MME | MME | MMB | MMB | SEED(-I) | MMMU | MMMU | VQA | GQA | SQA | POPE |
Bunny-695K | Frozen | 384 | 1402.0 | 286.1 | 73.1 | 72.3 | 64.6(71.9) | 40.7 | 39.3 | 80.6 | 62.4 | 75.3 | 85.8 |
1152 | 1473.1 | 332.1 | 72.8 | 72.0 | 64.1(71.3) | 41.8 | 38.3 | 80.6 | 62.4 | 75.6 | 86.6 | ||
Bunny-695K + LLaVA-665K | Trainable | 384 | 1490.5 | 316.4 | 72.4 | 72.0 | 63.7(71.3) | 40.0 | 38.3 | 81.3 | 63.0 | 74.4 | 86.4 |
1152 | 1503.5 | 336.4 | 74.4 | 74.3 | 64.2(71.4) | 38.8 | 38.6 | 81.6 | 63.1 | 75.0 | 87.0 |
Higher Resolution.
The ability of understanding high-resolution images and capturing the fine-grained details is crucial for a helpful MLLM. Given that the default image size of SigLIP-SO is 384384, we try two approaches to scale up image resolution: positional embedding interpolation and sub-image decomposition. Through positional embedding interpolation, the vision encoder is adapted to the new resolution during fine-tuning, i.e. 448448, and an image is represented with 1024 tokens, resulting a 20% training time increase. Tab. 9 shows that the improvement is marginal. For sub-image decomposition, we leverage S2-Wrapper [65], a simple mechanism that extends pretrained vision model to multiple image scales, which processes images at multiple scales by interpolating, splitting, encoding, merging, and concatenating features, resulting in a multi-scale representation with 3 times channel dimension while maintaining the token length. It would increase the training time for 23%. As shown in Tab. 10, S2-Wrapper leads to a huge performance improvement while using Bunny-695K + LLaVA-665K. Considering the performance and high resolution, we take S2-Wrapper as the final solution.
LLM | S2 | Fine-tuning Data | VE | MME | MME | MMB | MMB | SEED(-I) | MMMU | MMMU | VQA | GQA | SQA | POPE |
Phi-3-Mini | wo | B-695K | F | 1402.0 | 286.1 | 73.1 | 72.3 | 64.6(71.9) | 40.7 | 39.3 | 80.6 | 62.4 | 75.3 | 85.8 |
B-695K + L-665K | T | 1490.5 | 316.4 | 72.4 | 72.0 | 63.7(71.3) | 40.0 | 38.3 | 81.3 | 63.0 | 74.4 | 86.4 | ||
3:7 Averaging | 1495.2 | 338.9 | 74.0 | 73.5 | 64.5(72.1) | 40.1 | 39.1 | 81.5 | 63.5 | 75.2 | 86.7 | |||
Phi-3-Mini | w | B-695K | F | 1473.1 | 332.1 | 72.8 | 72.0 | 64.1(71.3) | 41.8 | 38.3 | 80.6 | 62.4 | 75.6 | 86.6 |
B-695K + L-665K | T | 1503.5 | 336.4 | 74.4 | 74.3 | 64.2(71.4) | 38.8 | 38.6 | 81.6 | 63.1 | 75.0 | 87.0 | ||
3:7 Averaging | 1503.9 | 362.9 | 74.1 | 74.1 | 64.6(71.7) | 40.2 | 38.8 | 81.7 | 63.4 | 76.3 | 87.0 | |||
Phi-3-Mini | w | B-695K + A-Ins | F | 1552.4 | 306.8 | 74.0 | 73.8 | 62.8(71.0) | 40.6 | 38.1 | 81.0 | 62.3 | 75.0 | 86.7 |
B-695K + L-665K + A-Ins | T | 1590.1 | 343.9 | 75.6 | 74.4 | 64.4(71.9) | 42.4 | 38.1 | 82.0 | 62.9 | 77.8 | 86.9 | ||
3:7 Averaging (Bunny-4B) | 1581.5 | 361.1 | 75.7 | 74.2 | 64.9(72.5) | 41.4 | 38.4 | 82.1 | 63.2 | 78.3 | 87.2 | |||
Llama-3-8B | wo | B-695K | F | 1542.7 | 344.6 | 75.7 | 74.8 | 65.2(73.1) | 42.4 | 37.1 | 82.0 | 64.4 | 78.7 | 85.9 |
B-695K + L-665K | T | 1562.2 | 319.6 | 75.3 | 74.4 | 65.2(72.5) | 43.4 | 38.0 | 82.4 | 64.1 | 77.4 | 86.4 | ||
5:5 Averaging | 1588.9 | 321.1 | 77.2 | 76.7 | 65.9(73.3) | 42.8 | 39.0 | 82.6 | 64.8 | 80.4 | 86.9 | |||
Llama-3-8B | w | B-695K | F | 1614.7 | 310.0 | 75.6 | 75.4 | 65.2(72.6) | 42.3 | 38.2 | 81.5 | 63.4 | 80.4 | 87.3 |
B-695K + L-665K | T | 1571.2 | 314.3 | 75.9 | 75.5 | 64.8(72.6) | 42.2 | 37.6 | 82.4 | 64.8 | 79.2 | 86.8 | ||
6:4 Averaging | 1607.8 | 324.6 | 76.5 | 76.3 | 65.9(73.6) | 43.7 | 38.8 | 82.3 | 64.5 | 80.4 | 87.4 | |||
Llama-3-8B | w | B-695K + A-Ins | F | 1599.6 | 348.9 | 76.3 | 75.6 | 66.2(73.4) | 41.6 | 38.3 | 81.9 | 64.1 | 79.5 | 86.4 |
B-695K + L-665K + A-Ins | T | 1649.7 | 341.1 | 77.2 | 76.5 | 65.5(73.1) | 40.4 | 38.8 | 82.8 | 64.1 | 79.0 | 87.3 | ||
4.2:5.8 Averaging (Bunny-8B) | 1644.1 | 367.5 | 78.1 | 77.2 | 66.2(73.5) | 43.3 | 39.0 | 82.9 | 64.0 | 79.9 | 87.2 |
Weight Merging.
During ablation studies, we obtain numerous models with the same architecture under different settings. We explore the potential benefits of merging two models by weighted averaging. As shown in Tab. 11, simple linear merging greatly enhances performance, with the best results forming our Bunny-4B/8B models. This method is more efficient and effective than hyper-parameter searching for improving model performance. We also observe that the results on some benchmarks (e.g. MME) are very sensitive to the weight factor.
Vision Encoder | LLM | MME | MME | MMB | MMB | SEED(-I) | MMMU | MMMU | VQA | GQA | SQA | POPE |
EVA02-CLIP-L (0.4B) | Phi-1.5 (1.3B) | 1213.7 | 278.9 | 60.9 | 56.8 | 56.4(64.1) | 30.0 | 28.4 | 76.5 | 60.4 | 58.2 | 86.1 |
StableLM-2 (1.6B) | 1301.0 | 235.0 | 58.4 | 56.4 | 55.3(62.8) | 29.8 | 29.4 | 74.6 | 56.7 | 60.0 | 84.8 | |
Phi-2 (2.7B) | 1421.0 | 285.4 | 68.6 | 67.4 | 62.2(70.2) | 35.9 | 32.6 | 78.9 | 62.3 | 69.1 | 87.1 | |
SigLIP-SO (0.4B) | Phi-1.5 (1.3B) | 1230.0 | 237.5 | 61.2 | 59.7 | 57.7(65.3) | 30.0 | 29.1 | 78.0 | 61.1 | 61.3 | 85.8 |
StableLM-2 (1.6B) | 1366.8 | 236.1 | 65.1 | 62.8 | 58.8(67.5) | 29.9 | 29.8 | 78.9 | 60.9 | 61.1 | 85.9 | |
Phi-2 (2.7B) | 1488.8 | 289.3 | 69.2 | 68.6 | 62.5(70.7) | 38.2 | 33.0 | 79.8 | 62.5 | 70.9 | 86.8 | |
Phi-3-Mini (3.8B) | 1581.5 | 361.1 | 75.7 | 74.2 | 64.9(72.5) | 41.4 | 38.4 | 82.1 | 63.2 | 78.3 | 87.2 | |
Llama-3-8B | 1644.1 | 367.5 | 78.1 | 77.2 | 66.2(73.5) | 43.3 | 39.0 | 82.9 | 64.0 | 79.9 | 87.2 |